OpenStudy (anonymous):

Find the general solution to these differential equations. xy+y'=5x

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

answer was y=5+Ce^(-x^2/2)

OpenStudy (anonymous):

woudln't it be \[-Ce^\frac{ x^2 }{ 2 } + 5\]

OpenStudy (anonymous):

@agent0smith

zepdrix (zepdrix):

"wouldn't it be.." No it wouldn't. Did we make a small error near the end somewhere? Your answer looks awfully close to the book answer.

zepdrix (zepdrix):

Need to see some steps? :o

OpenStudy (sirm3d):

i got \[y=5+Ce^{-\tfrac{1}{2}x^2}\]

terenzreignz (terenzreignz):

Integrating factors? Starting with... \[\huge \frac{dy}{dx}+xy=5x\]

terenzreignz (terenzreignz):

Integrating factor is \[\large \int x dx=\frac12x^2\]So multiply \[\huge e^{\frac{x^2}{2}}\]to the entire expression... \[\huge e^{\frac{x^2}{2}}\frac{dy}{dx}+ e^{\frac{x^2}{2}}xy=5x e^{\frac{x^2}{2}}\]

terenzreignz (terenzreignz):

Turns out the left-hand side is just \[\huge \frac{d}{dx}\left(y e^{\frac{x^2}{2}}\right)=5x e^{\frac{x^2}{2}}\]